# Passive defect driven morphogenesis in nematic membranes

### v1

@@ Abstract
Finally, this study reveals a novel mechanism for the spontaneous fusion of half-integer topological defects on fluid membranes. Consider the special case $\sigma=0$ and $k_{1}\!=\!k_{2}\!=\!k_{3}\!=\!k/3$. It is known that the energy minimizing nematic field with a total charge of $+1$ on a flat disc has two $+1/2$ defects. In this case, the energy scales, up to numerical pre-factors of order $1$, as ${\cal F}_{f}\sim\pi k/3\log(R/\Delta)$ [44]. However, the free-energy of a conical surface with an aster at the apex scales sub-linearly with $k$, (i.e ${\cal F}_{c}\sim\pi k_{B}\sqrt{(4k/3k_{B}-1)})\log(R/\Delta)$). Therefore, a critical threshold $(k/3k_{B})_{c}=2+\sqrt{3}$ arises from the balance between the energies of these two states. Thus, if $(k/3k_{B})>(k/3k_{B})_{c}$, the pair of $+1/2$ topological defects can spontaneously fuse by deforming the surface out of plane. Indeed, past works have shown that this process can occur for prescribed conical surfaces [25].
Next we discuss the relevance of this mechanism in biological systems, see Sec. 5 in [42]. For suspensions of cytoskeletal filaments, the Frank constant ranges from $k/h\sim 0.1-1~{}p$N for actin to $100~{}p$N for microtubules, [45, 46, 47, 48], where $h=0.1$ $\mu$m is the average thickness of a cytoskeletal layer [49, 50, 51]. In the bending dominated regime ($k_{B}\gg\sigma R^{2}$), taking the bending modulus of a vesicle $k_{B}\sim 10-100~{}pN\cdot nm$ [52, 53, 54], we obtain that the ratio $k/k_{B}\sim 1$ for actin and it is $k/k_{B}\sim 10^{2}-10^{3}$ for microtubules. In the tension dominated regime $k_{B}\ll\sigma R^{2}$, the threshold is controlled in addition by the system geometry via $R$. Taking the tension of a vesicle to be $10^{-3}$ N/m [52, 53, 54], we obtain that the critical lengthscale $R_{c}=\sqrt{k/\sigma}$ is comparable to the thickness $h$ and therefore negligible. These estimations suggest that in low-tension regimes, biological system can induce out-of-plane deformation via the mechanism described above. In fact, this phenomenon may have been observed in Ref. [37] that studied a suspension of microtubules encapsulated on vesicles. In this work, the authors report that lowering the surface tension leads to a spindle-like vesicles with two $+1$ defects localized at the spindle poles. In addition, the equilibrium deformation by topological defects can also influence shape dynamics in systems driven by out-of-equilibrium processes [37, 36, 55]. For instance, because the aster tend to facilitate surface deformations, it can be a nucleation point for out-of-plane deformations on fluid membranes, bacterial biofilms or cell monolayers.
@@ Acknowledgements.
We are grateful to Isabelle Bonnet, Mathieu Dedenon, Karsten Kruse, Jean François Joanny, Jacques Prost, and Feng-Ching Tsai for insightful discussions. DJGP acknowledges funding from the Swiss National Science Foundation under starting grant TMSGI2 211367. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 847718. Q.C acknowledges funding from Institut Curie EuReCa PhD Programme. This publication reflects only the author’s view and that the European Research Agency is not responsible for any use that may be made of the information it contains. References cited in SI: [56, 57, 58, 59]

### v2

@@ Abstract
Finally, this study reveals a novel mechanism for the spontaneous fusion of half-integer topological defects on fluid membranes. Consider the special case $\sigma=0$ and $k_{1}\!=\!k_{2}\!=\!k_{3}\!=\!k/3$. It is known that the energy minimizing nematic field with a total charge of $+1$ on a flat disc has two $+1/2$ defects. In this case, the energy scales, up to numerical pre-factors of order $1$, as ${\cal F}_{f}\sim\pi k/3\log(1/\Delta)$ [46]. However, the free-energy of a conical surface with an aster at the apex scales sub-linearly with $k$, (i.e ${\cal F}_{c}\sim\pi k_{B}\sqrt{(4k/3k_{B}-1)})\log(1/\Delta)$). Therefore, a critical threshold $(k/3k_{B})_{c}={\cal O}(1)$ arises from the balance between the energies of these two states. Thus, if $(k/3k_{B})>(k/3k_{B})_{c}$, the pair of $+1/2$ topological defects can spontaneously fuse by deforming the surface out of plane. Indeed, past works have shown that this process can occur for prescribed conical surfaces [25].
This mechanism for spontaneous deformation of fluid membranes can drive out-of-plane deformations of biological system, such as a film of cytoskeletal filaments on a supported lipid bilayer or a cell monolayer on a supported elastic substrate, see Sec. 6 in [44]. For instance, actin and microtubule films can exhibit nematic phases [47, 48, 49]. The measured reduced Frank constant is in the range of $k=10^{-8}-10^{-7}~{}\mu N\mu m$ for actin filament films [50, 51] and is $k=10^{-3}~{}\mu N\mu m$ for microtubule films [52]. Using typical values for the bending rigidity and surface tension of lipid bilayers and ignoring other possible elastic contributions, allows one to evaluate the proximity to the instability threshold (S5). In particular, in the bending dominate regime $k_{B}\gg\sigma R^{2}$, the critical threshold is of order of $(k/k_{B})_{c}\sim 1$, and the ratio is $k/k_{B}=1$ for a film of actin filaments and is $k/k_{B}=10^{4}-10^{5}$ for a film of microtubules. In the tension dominated regime $k_{B}\ll\sigma R^{2}$, the critical threshold is $(k/\sigma R^{2})_{c}\sim 1$ and therefore it can vary with the disc radius $R$. One can determine an upper bound by replacing the radius $R$ with the layer thickness $h$, which is the smallest lengthscale. This bound is $k/\sigma h^{2}<10$ for a film of both types of cytoskeletal filaments. These estimations suggest that in low-tension regimes, biological system can induce out-of-plane deformation by relieving elastic stresses at integer topological defects. In fact, this phenomenon may have been observed in Ref. [38] that studied a film of microtubules encapsulated on a vesicle. In this work, the authors reported that lowering the surface tension leads to spindle-like vesicles with two $+1$ defects localized at the spindle poles.
Shape dynamics driven by out-of-equilibrium processes could also be influenced by this mechanism. For instance, because the aster tends to facilitate surface deformations, it can be a nucleation point for out-of-plane deformations on active fluid membranes, like bacterial biofilms or cell monolayers [38, 37, 53, 40, 29].
@@ Acknowledgements.
We are grateful to Isabelle Bonnet, Mathieu Dedenon, Karsten Kruse, Jean François Joanny, Jacques Prost, and Feng-Ching Tsai for insightful discussions. DJGP acknowledges funding from the Swiss National Science Foundation under starting grant TMSGI2 211367. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 847718. Q.C acknowledges funding from Institut Curie EuReCa PhD Programme. This publication reflects only the author’s view and that the European Research Agency is not responsible for any use that may be made of the information it contains.
References cited in SI: [38, 54, 37, 39, 40, 55, 29, 47, 48, 49, 56, 57, 58, 59, 50, 51, 60, 61, 62, 52, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 46, 75, 76]